find the total cost of a pair of jeans that cost $32 and has an 8% sales tax
Found 2 solutions by neatmath, fractalier:
Answer by neatmath(302)

Total Cost=cost of item + cost of item*sales tax %

T=C+C*S

T=32+32*8%

T=32+2.56

T=34.56

Thus, the total cost for the jeans (after sales tax) would be $34.56

I hope this helps!

Answer by fractalier(6550)

If the tax on $32 is 8%, we can find the tax by multiplying and add that to the $32.
So 8% = .08
.08 x $32 = $2.56
and the total cost is
$34.56